# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ 4.3.4 Electrocardiogram (ECG) Classification
Disorders pertaining to the heart or blood vessels are collectively referred to as Cardiovascular Diseases (CVD) [233]. According to the American Heart Association’s 2023 statistics, CVD has emerged as the leading cause of death worldwide. In 2020, 19.05 million deaths were recorded from CVD globally, which signifies an increase of 18.71% from 2010, and it is believed that this number will rise to 23.6 million by 2030 [234]. Blood clots and vascular blockages caused by CVDs can cause myocardial infarction, stroke or even death [233]. Generally, early diagnosis has been shown to reduce the mortality rate of CVDs, and Electrocardiogram (ECG) signals play a crucial role in diagnosing various cardiac abnormalities and monitoring heart health. However, ECG signal has characteristics of high noise and high complexity, making it time-consuming and labor-intensive to identify certain diseases using traditional methods. The traditional approach is tedious and requires the expertise of a medical specialist. Over the past decades, the task of Long-term ECG recording classification has been significantly facilitated for cardiologists through the adoption of computerized ECG recognition practices. Throughout this period, feature extraction methods have predominantly relied on manual techniques, encompassing diverse approaches such as wave shape functions [235], wavelet-based features [236], ECG morphology [237], hermite polynomials [238], and Karhunen-Loeve expansion of ECG morphology [239], among others. These extracted features are subsequently subjected to classification using various machine learning algorithms.
@@
More recently, the advent of deep learning has revolutionized the field by enabling automatic feature learning directly from ECG signals. This advancement holds significant promise in the realm of automated ECG classification, offering clinicians a tool for swift and accurate diagnosis. Based on this, several deep learning architectures have been proposed for feature learning of ECG signals. For instance, Acharya et al. [240] developed a 9-layer CNN model to automatically identify five categories of heartbeats in ECG signals. A similar model was also developed in [241], However, ECG signals often vary significantly in length, as they may contain different numbers of heartbeats. CNNs typically require fixed-length inputs, which may necessitate preprocessing steps such as padding or truncation, potentially losing important temporal information. For this reason, several architectures have leveraged RNN in ECG classification, as seen in [242], [243] and [244], among others. While RNNs are capable of handling sequential data, they also have limitations in capturing local patterns or short-term dependencies effectively. In ECG signals, local features such as specific waveforms or intervals can be crucial for classification. For this reason, recent works have proposed hybrid models which combine the strengths of both CNNs and RNNs to overcome some of these limitations [245].
@@
The work of Rai et al. [246] developed a hybrid CNN-LSTM network to evaluate the optimum performing model for myocardial infarction detection using ECG signals. The authors then experimented on 123,998 ECG beats obtained from the PTB diagnostic database (PTBDB) and MIT-BIH arrhythmia database (MITDB), and the result showed that by combining the capabilities of both CNN and LSTM, improved classification accuracy can be achieved. Also, in [247], a CNN architecture was developed to extract morphological features from ECG signals. For the purpose of determining the degree of heart rate variability, another composite structure was designed using LSTM and a collection of manually created statistical features. Following that, a hybrid CNN-LSTM architecture is built using the two independent biomarkers to classify cardiovascular artery diseases, and experiments were carried out on two distinct datasets. The first is a partly noisy in-house dataset collected using an inexpensive ECG sensor, and the other is a corpus taken from the MIMIC II waveform dataset. The hybrid model proposed in the work achieved an overall classification accuracy of 88% and 93%, respectively, which surpasses the performance of standalone architectures.
@@
An automated diagnosis method based on Deep CNN and LSTM architecture was presented in [248] to identify Congestive Heart Failure (CHF) from ECG signals. Specifically, CNN was used to extract deep features, and LSTM was employed to exploit the extracted features to achieve the CHF detection goal. The model was tested using real-time ECG signal datasets, and the results showed that the AUC was 99.9%, the sensitivity was 99.31%, the specificity was 99.28%, the F-Score was 98.94%, and the accuracy was 99.52%. However, since ECG signals can vary in length due to differences in recording durations or patient conditions. LSTMs are capable of handling variable-length sequences, but traditional CNNs typically require fixed-length inputs. Therefore, fusing these features effectively in a hybrid model can be challenging. Also, Hybrid CNN-RNN models can be computationally intensive, especially when processing long ECG sequences or large datasets. For this reason, recent research works have proposed the use of attention mechanisms to reduce the computational burden by enabling the model to selectively attend to informative features, focusing computational resources where they are most needed. Likewise, attention mechanisms can enable the model to attend to informative segments of the ECG signal, regardless of their length, allowing for more flexible processing of variable-length sequences.
@@
Several researchers have leveraged attention mechanisms in standalone and hybrid architectures for improved performance. For instance, in the work of Chun-Yen et al. [249], CNN layers were used to extract main features, while LSTM and attention were included to enhance the model’s feature learning capabilities. Experiments on a 12-lead KMUH ECG dataset showed that the model had high recognition rates in classifying normal and abnormal ECG signals, compared to hybrid models without attention mechanisms. Wang et al. [250] presented a 33-layer CNN architecture with non-local convolutional block attention module (NCBAM). To extract the spatial and channel information, preprocessed ECG signals were first fed into the CNN architecture. A non-local attention further captured long-range dependencies of representative features along spatial and channel axes. Similarly, a spatio-temporal attention-based convolutional recurrent neural network (STA-CRNN) was presented in [251] with the aim of concentrating on representative features in both the spatial and temporal dimensions. The CNN subnetwork, spatiotemporal attention modules, and RNN subnetwork made up the STA-CRNN and according to findings, the STA-CRNN model was able to classify eight different forms of arrhythmias and normal rhythm with an average F1 score of 0.835.
@@
Combining hybrid deep learning models with attention mechanisms for ECG feature learning is a promising approach that has already shown potential in ECG feature learning, according to reviewed literature. Future research can further explore semi-supervised and self-supervised learning techniques to leverage large amounts of unlabeled ECG data. This could involve pre-training models on large-scale unlabeled datasets using self-supervised learning objectives. Also, deep learning models have been leveraged in the generation of synthetic ECG signals to augment real signals, as seen in [252] where a GAN model was developed to generate ECG signals that correspond with available clinical data. The GAN model used two layers of BiLSTM networks for the generator and CNN for the discriminator, and trained using the 48 ECG recordings of different users from the MIT-BIH dataset. The authors then compared their model with a Recurrent neural network autoencoder (RNN-AE) model and a recurrent neural network variational autoencoder (RNN-VAE) model, and the results showed that their model exhibited the fastest convergence of its loss function to zero.
@@ 4.3.5 Electroencephalography (EEG) Classification
Three-dimensional scalp surface electrode readings provide a dynamic time series that is called Electroencephalogram (EEG) signal [253]. Brain waves obtained from an EEG can effectively depict both the psychological and pathological states of a human. The human brain is acknowledged to be a fascinating and incredibly complicated structure. Numerous brain signals, including functional magnetic resonance imaging (fMRI), near-infrared spectroscopy (NIRS), electroencephalograms (EEGs), and functional near-infrared spectroscopy (fNIR), among others have been collected and used to study the brain [254]. Due to the EEG’s non-invasive, affordable, accessible, and excellent temporal resolution characteristics, it has become the most utilised approach. However, the signal-to-noise ratio of EEG signal is low, meaning that sources with no task-relevant information frequently have a stronger effect on the EEG signal than those that do. These characteristics often make end-to-end feature learning for EEG data substantially more challenging [253]. Based on this, several methods have been leveraged for improved feature extraction in EEG signals across several domains including Motor imagery [255], anxiety disorder [256], epileptic seizure detection [257], sleep pattern analysis and disorder detection [258, 259], and Alzheimer’s disease detection [260], and many others.
@@
EEG Motor Imagery (MI) is a technique used to study brain activity associated with the imagination of movement. It involves recording electrical activity generated by the brain through electrodes placed on the scalp. MI tasks typically involve imagining performing a specific motor action, such as moving a hand or foot, without physically executing the movement, and has been leveraged in smart healthcare applications such as post-stroke rehabilitation and mobile assistive robots, among others [261]. Prior to the advent of deep learning, motor imagery EEG data are passed through various steps before classification using traditional ML techniques. Pre-processing, feature extraction, and classification are the three primary stages that traditional approaches usually take while processing MI-EEG signals. Pre-processing includes a number of operations, including signal filtering (choosing the most valuable frequency range for MI tasks), channel selection (identifying the most valuable EEG channels for MI tasks), signal normalisation (normalising each EEG channel around the time axis), and artefact removal (removing noise from MI-EEG signals). Independent component analysis (ICA) is the most often utilised technique for removing artefacts [262, 263, 264]. In contrast to the traditional approach, deep learning architectures can automatically extract complex features from raw MI-EEG data without the need for laborious feature extraction and pre-processing. Based on this, several deep learning architectures have been proposed for MI-EEG feature learning, as seen in [265], [266] and [267], among others. For instance [253], categorized MI-EEG signals using three CNNs with varying architectures, and the number of convolutional layers varied from two layers to a five-layer deep ConvNet to a thirty-one-layer residual network.
@@
In Dai et al. [268], the authors proposed an approach for classifying MI-EEG signals which blend variational autoencoder with CNN architecture. The VAE decoder was used to fit the Gaussian distribution of EEG signals, and the time, frequency, and channel information from the EEG signal were combined to create a novel representation of input, and the proposed CNN-VAE method was optimised for the input. Experiments showed that by combining both deep learning architectures, improved features were learnt, which led to a high classification performance on the BCI Competition IV dataset 2b. Li et al. [269] employed optimal wavelet packet transform (OWPT) for the generation of feature vectors from MI-EEG signals. These vectors were then utilized to train an LSTM network which demonstrated satisfactory performance on dataset III from the BCI Competition 2003. However, the model has excessively intricate structure. To address this, Feng et al. [270] introduced a technique that merges continuous wavelet transform (CWT) with a simplified convolutional neural network to enhance the accuracy of recognizing MI-EEG signals. By employing CWT, MI-EEG signals were converted into time-frequency image representations. Subsequently, these image representations were fed into the SCNN for feature extraction and classification. Experiments on the BCI Competition IV Dataset 2b demonstrate that, on average, the classification accuracy across nine subjects reached 83.2%. However, the computational complexity of the model was quite high, due to the processing of time-frequency image representations. The conversion of MI-EEG signals into time-frequency images using CWT requires significant computational resources.
@@
The authors in [271] introduced a classification framework based on Long Short-Term Memory (LSTM) to improve the accuracy of classifying four-class motor imagery signals from EEG. The authors sliding window technique to capture time-varying EEG signal data, and employed an overlapping-band-based Filter Bank Common Spatial Patterns (FBCSP) method to extract subject-specific spatial features. Experiments on the BCI Competition IV dataset 2a, showed that their model achieved an average accuracy of 97%, compared to existing methods. Also, in the classification of Alzheimer’s disease, Zhao et al. [272] employed a deep learning network to analyse EEG data. The deep learning model was evaluated on a dataset that consist of fifteen (15) patients with clinically confirmed Alzheimer’s disease and fifteen (15) healthy individuals, and results showed that improved features were learnt and compared the results to the traditional methods. This has prompted the use of deep learning in Alzheimer’s disease detection, as seen in [273], where the authors used CNN for diagnosing Alzheimer’s Disease. To address challenges posed by limited data and overfitting in deep learning models designed for Alzheimer’s Disease detection, the authors explored the use of overlapping sliding windows to augment the EEG data collected from 100 subjects (comprising 49 AD patients, 37 mild cognitive impairment patients, and 14 healthy controls subjects). After assembling the augmented dataset, a modified Deep Pyramid Convolutional Neural Network (DPCNN) was used to classify the enhanced EEG signals. In epilepsy detection, [274] developed three deep learning architectures (CNN, LSTM, and hybrid CNN-LSTM), with each model chosen for its effectiveness in handling the intricate characteristics of EEG data. Each architecture offers distinct advantages, with CNN excelling in spatial feature extraction, LSTM in capturing temporal dynamics, and the hybrid model combining these strengths. The CNN model, consisting of 31 layers, attained the highest accuracy, achieving 91% on the first benchmark dataset and 82% on the second dataset using a 30-second threshold, selected for its clinical significance.

### v6

@@ 4.3.4 Electrocardiogram (ECG) Classification
Disorders pertaining to the heart or blood vessels are collectively referred to as Cardiovascular Diseases (CVD) [283]. According to the American Heart Association’s 2023 statistics, CVD has emerged as the leading cause of death worldwide. In 2020, 19.05 million deaths were recorded from CVD globally, which signifies an increase of 18.71% from 2010, and it is believed that this number will rise to 23.6 million by 2030 [284]. Blood clots and vascular blockages caused by CVDs can cause myocardial infarction, stroke, or even death [283]. Generally, early diagnosis has been shown to reduce the mortality rate of CVDs, and Electrocardiogram (ECG) signals play a crucial role in diagnosing various cardiac abnormalities and monitoring heart health. However, ECG signal has characteristics of high noise and high complexity, making it time-consuming and labor-intensive to identify certain diseases using traditional methods. The traditional approach is tedious and requires the expertise of a medical specialist. Over the past decades, the task of Long-term ECG recording classification has been significantly facilitated for cardiologists through the adoption of computerized ECG recognition practices. Throughout this period, feature extraction methods have predominantly relied on manual techniques, encompassing diverse approaches such as wave shape functions [285], wavelet-based features [286], ECG morphology [287], hermite polynomials [288], and Karhunen-Loeve expansion of ECG morphology [289], among others. These extracted features are subsequently subjected to classification using various machine learning algorithms.
@@
More recently, the advent of deep learning has revolutionized the field by enabling automatic feature learning directly from ECG signals. This advancement holds significant promise in the realm of automated ECG classification, offering clinicians a tool for swift and accurate diagnosis. Based on this, several deep learning architectures have been proposed for feature learning of ECG signals. For instance, Acharya et al. [290] developed a 9-layer CNN model to automatically identify five categories of heartbeats in ECG signals. A similar model was also developed in [291]. However, ECG signals often vary significantly in length, as they may contain different numbers of heartbeats. CNNs typically require fixed-length inputs, which may necessitate preprocessing steps such as padding or truncation, potentially losing important temporal information. For this reason, several architectures have leveraged RNN in ECG classification, as seen in [292], [293] and [294], among others. While RNNs are capable of handling sequential data, they also have limitations in capturing local patterns or short-term dependencies effectively. In ECG signals, local features such as specific waveforms or intervals can be crucial for classification. For this reason, recent works have proposed hybrid models which combine the strengths of both CNNs and RNNs to overcome some of these limitations [295].
@@
Hybrid Models: The work of Rai et al. [296] developed a hybrid CNN-LSTM network to evaluate the optimum performing model for myocardial infarction detection using ECG signals. The authors then experimented on 123,998 ECG beats obtained from the PTB diagnostic database (PTBDB) and MIT-BIH arrhythmia database (MITDB), and the result showed that by combining the capabilities of both CNN and LSTM, improved classification accuracy can be achieved. Also, in [297], a CNN architecture was developed to extract morphological features from ECG signals. For the purpose of determining the degree of heart rate variability, another composite structure was designed using LSTM and a collection of manually created statistical features. Following that, a hybrid CNN-LSTM architecture is built using the two independent biomarkers to classify cardiovascular artery diseases, and experiments were carried out on two distinct datasets. The first is a partly noisy in-house dataset collected using an inexpensive ECG sensor, and the other is a corpus taken from the MIMIC II waveform dataset. The hybrid model proposed in the work achieved an overall classification accuracy of 88% and 93%, respectively, which surpasses the performance of standalone architectures.
@@
An automated diagnosis method based on Deep CNN and LSTM architecture was presented in [298] to identify Congestive Heart Failure (CHF) from ECG signals. Specifically, CNN was used to extract deep features, and LSTM was employed to exploit the extracted features to achieve the CHF detection goal. The model was tested using real-time ECG signal datasets, and the results show that the AUC was 99.9%, the sensitivity was 99.31%, the specificity was 99.28%, the F-Score was 98.94%, and the accuracy was 99.52%. Similar work is reported in [299] whereby CNN is combined with LSTM for ECG classification. The hybrid model consists of three blocks of convolutional and maximum pooling layers with dropout followed by an LSTM with drop. Furthermore, Shapley is utilized to determine the contribution of each ECG sample on the prediction to improve model interpretability. However, since ECG signals can vary in length due to differences in recording durations or patient conditions. LSTMs are capable of handling variable-length sequences, but traditional CNNs typically require fixed-length inputs. Therefore, fusing these features effectively in a hybrid model can be challenging. Furthermore, Hybrid CNN-RNN models can be computationally intensive, especially when processing long ECG sequences or large datasets. For this reason, recent research works have proposed the use of attention mechanisms to reduce the computational burden by enabling the model to selectively attend to informative features, focusing computational resources where they are most needed. Likewise, attention mechanisms can enable the model to attend to informative segments of the ECG signal, regardless of their length, allowing for more flexible processing of variable-length sequences.
@@
Several researchers have leveraged attention mechanisms in standalone and hybrid architectures for improved performance. For instance, in the work of Chun-Yen et al. [300], CNN layers were used to extract main features, while LSTM and attention were included to enhance the model’s feature learning capabilities. Experiments on a 12-lead KMUH ECG dataset showed that the model had high recognition rates in classifying normal and abnormal ECG signals, compared to hybrid models without attention mechanisms. Wang et al. [301] presented a 33-layer CNN architecture with non-local convolutional block attention module (NCBAM). To extract the spatial and channel information, preprocessed ECG signals were first fed into the CNN architecture. A non-local attention further captured long-range dependencies of representative features along spatial and channel axes. Similarly, a spatio-temporal attention-based convolutional recurrent neural network (STA-CRNN) was presented in [302] with the aim of concentrating on representative features in both the spatial and temporal dimensions. The CNN subnetwork, spatiotemporal attention modules, and RNN subnetwork formed the STA-CRNN and according to findings, the STA-CRNN model was able to classify eight different forms of arrhythmias and normal rhythm with an average F1 score of 0.835. In [303], the SE module is utilized in between convolutional layers and LSTM. This approach allows the model to focus on relevant channels before temporal dependencies are captured by the LSTM, improving the classification performance.
In [304], a guided spatial attention mechanism is introduced to incorporate domain knowledge into the model, enhancing ECG classification performance. The guided spatial attention mechanism has an autoencoder-like structure, where the feature maps are downsampled by the encoder and upsampled by the decoder. Additionally, skip connections are employed to improve information flow between the encoder and decoder. Furthermore, an attention loss term, based on the attention weights is introduced to jointly train the guided attention mechanisms. In [305], ECG signals are classified into sleep stages using deep learning. The approach is divided into three stages. In the first stage, a deep learning model consisting of convolutional layers and liquid time-constant network [306] is employed to compute kurtosis and skewness of the ECG signals. In the second stage, a deep learning model with convolutional layers and SE attention is used to enhance the prediction of the minority sleep stage class. In the final stage, the outputs of the models are concatenated and classified into sleep stage categories. Combining hybrid deep learning models with attention mechanisms for ECG feature learning is a promising approach that has already shown potential in ECG feature learning, according to reviewed literature. Future research can further explore semi-supervised and self-supervised learning techniques to leverage large amounts of unlabeled ECG data. This could involve pre-training models on large-scale unlabeled datasets using self-supervised learning objectives.
@@
Generative Models: Deep learning models have been leveraged in the generation of synthetic ECG signals to augment real signals, as seen in [307] where a GAN model was developed to generate ECG signals that correspond with available clinical data. The GAN model used two layers of BiLSTM for the generator and CNN for the discriminator, and trained using the 48 ECG recordings of different users from the MIT-BIH dataset. The authors then compared their model with a Recurrent neural network autoencoder (RNN-AE) model and a recurrent neural network variational autoencoder (RNN-VAE) model, and the results show that their model exhibited the fastest convergence of its loss function to zero. Similar work is reported in [308] in which GAN and autoencoder are employed to generate ECG signals to address the issue of imbalanced dataset as shown in Figure 32. The generator and encoder create synthetic and real low-dimensional ECG samples, respectively. The discriminator classifies the synthetic and real low-dimensional ECG samples as real or fake, while the decoder reconstructs the ECG signals from the synthetic low-dimensional samples. This approach further refines data transformation and reconstruction quality, generating more realistic ECG signals. In [309], cluster-based GAN is employed to improve the diversity and quality of the generated synthetic ECG signals. $K$-means algorithm is first employed to partition the ECG signals into $k$ distinct groups. Then, the GAN is trained using the cluster centroids as initial reference points to guide the GAN in generating synthetic signals. Table 12 lists the summary of state-of-the-art ECG classification.
Figure 32: The architecture of CECG-GAN [308].
Table 12: Summary of state-of-the-art ECG classification.
Ref. Description Results (Datasets and Metrics)
[303], 2024 A hybrid model consists of three blocks of convolutional layer, SE module and an LSTM MIT-BIH ECG
Accuracy: 98.5%
[299], 2024 A lightweight hybrid model consists of three blocks of convolutional and maximum pooling layers and an LSTM MIT-BIH and LTAF
Accuracy: 98.24%
[300], 2022 A hybrid model consists of three convolutional pipelines with different kernel size, two LSTM layers and an attention module KMUH
Accuracy: 96.02%
CPSC-2018
Accuracy: 94.05%
[304], 2024 This study proposed a guided spatial attention mechanism to enhance ECG classification performance CPSC2018 and PTB-XL Chapman Cinc2017
F1-score: 77.56% (STC), 88.53% (PC), 82.54% (WPW), 85.37% (AF)
[305], 2024 A novel three-stage approach for sleep stage classification extracts statistical features using a Feature Imitating Network, enhances minority sleep stage detection, and integrates outputs for five-class classification MIT-BIH Polysomnographic
Accuracy: 80.79%
[308], 2024 This study proposed a generative model by combining conditional GAN and autoencoder CSPC2020 dataset
F1-score: 0.9823 (Normal), 0.9824 (Premature ventricular), 0.9825 (Premature Supraventricular)
[309], 2024 This study introduces a generative model by combining GAN and K-means MIT-BIH arrhythmia
Accuracy: 99.66%
@@ 4.3.5 Electroencephalography (EEG) Classification
Three-dimensional scalp surface electrode readings provide a dynamic time series that is called Electroencephalogram (EEG) signal [310]. Brain waves obtained from an EEG can effectively depict both the psychological and pathological states of a human. The human brain is acknowledged to be a fascinating and incredibly complicated structure. Numerous brain signals, including functional magnetic resonance imaging (fMRI), near-infrared spectroscopy (NIRS), electroencephalograms (EEGs), and functional near-infrared spectroscopy (fNIR), among others have been collected and used to study the brain [311]. Due to the EEG’s non-invasive, affordable, accessible, and excellent temporal resolution characteristics, it has become the most utilized approach. However, the signal-to-noise ratio of EEG signal is low, meaning that sources with no task-relevant information frequently have a stronger effect on the EEG signal than those that do. These characteristics often make end-to-end feature learning for EEG data substantially more challenging [310]. Based on this, several methods have been leveraged for improved feature extraction in EEG signals across several domains including Motor imagery [312], anxiety disorder [313], epileptic seizure detection [314], sleep pattern analysis and disorder detection [315, 316], and Alzheimer’s disease detection [317], and many others.
@@
EEG Motor Imagery (MI) is a technique used to study brain activity associated with the imagination of movement. It involves recording electrical activity generated by the brain through electrodes placed on the scalp. MI tasks typically involve imagining performing a specific motor action, such as moving a hand or foot, without physically executing the movement, and has been leveraged in smart healthcare applications such as post-stroke rehabilitation and mobile assistive robots, among others [318]. Prior to the advent of deep learning, motor imagery EEG data are passed through various steps before classification using traditional ML techniques. Pre-processing, feature extraction, and classification are the three primary stages that traditional approaches usually take while processing MI-EEG signals. Pre-processing includes a number of operations, including signal filtering (choosing the most valuable frequency range for MI tasks), channel selection (identifying the most valuable EEG channels for MI tasks), signal normalization (normalizing each EEG channel around the time axis), and artefact removal (removing noise from MI-EEG signals). Independent component analysis (ICA) is the most often utilized technique for removing artefacts [319, 320, 321]. In contrast to the traditional approach, deep learning architectures can automatically extract complex features from raw MI-EEG data without the need for laborious feature extraction and pre-processing. Based on this, several deep learning architectures have been proposed for MI-EEG feature learning, as seen in [322], [323] and [324], among others. For instance [310], categorized MI-EEG signals using three CNNs with varying architectures, and the number of convolutional layers varied from two layers to a five-layer deep ConvNet to a thirty-one-layer residual network.
@@
In Dai et al. [325], the authors proposed an approach for classifying MI-EEG signals which blend variational autoencoder with CNN architecture. The VAE decoder was used to fit the Gaussian distribution of EEG signals, and the time, frequency, and channel information from the EEG signal were combined to create a novel representation of input, and the proposed CNN-VAE method was optimized for the input. Experiments showed that by combining both deep learning architectures, improved features were learnt, which led to a high classification performance on the BCI Competition IV dataset 2b. Li et al. [326] employed optimal wavelet packet transform (OWPT) for the generation of feature vectors from MI-EEG signals. These vectors were then utilized to train an LSTM network, which demonstrated satisfactory performance on dataset III from the BCI Competition 2003. However, the model has an excessively intricate structure. To address this, Feng et al. [327] introduced a technique that merges continuous wavelet transform (CWT) with a simplified convolutional neural network to enhance the accuracy of recognizing MI-EEG signals. By employing CWT, MI-EEG signals were converted into time-frequency image representations. Subsequently, these image representations were fed into the SCNN for feature extraction and classification. Experiments on the BCI Competition IV Dataset 2b demonstrate that, on average, the classification accuracy across nine subjects reached 83.2%. However, the computational complexity of the model was quite high, due to the processing of time-frequency image representations. The conversion of MI-EEG signals into time-frequency images using CWT requires significant computational resources.
@@
The authors in [328] introduced a classification framework based on Long Short-Term Memory (LSTM) to improve the accuracy of classifying four-class motor imagery signals from EEG. The authors used sliding window technique to capture time-varying EEG signal data, and employed an overlapping-band-based Filter Bank Common Spatial Patterns (FBCSP) method to extract subject-specific spatial features. Experiments on the BCI Competition IV dataset 2a, showed that their model achieved an average accuracy of 97%, compared to existing methods. Also, in the classification of Alzheimer’s disease, Zhao et al. [329] employed a deep learning network to analyze EEG data. The deep learning model was evaluated on a dataset that consisted of fifteen (15) patients with clinically confirmed Alzheimer’s disease and fifteen (15) healthy individuals, and results showed that improved features were learnt and compared the results to the traditional methods. This has prompted the use of deep learning in Alzheimer’s disease detection, as seen in [330], where the authors used CNN for diagnosing Alzheimer’s Disease. To address challenges posed by limited data and overfitting in deep learning models designed for Alzheimer’s Disease detection, the authors explored the use of overlapping sliding windows to augment the EEG data collected from 100 subjects (comprising 49 AD patients, 37 mild cognitive impairment patients, and 14 healthy controls subjects). After assembling the augmented dataset, a modified Deep Pyramid Convolutional Neural Network (DPCNN) was used to classify the enhanced EEG signals.